# Learning relevant contextual variables within Bayesian optimization

### v3

@@ Abstract
Contextual Bayesian Optimization (CBO) efficiently optimizes black-box functions with respect to design variables, while simultaneously integrating contextual information regarding the environment, such as experimental conditions. However, the relevance of contextual variables is not necessarily known beforehand. Moreover, contextual variables can sometimes be optimized themselves at additional cost, a setting overlooked by current CBO algorithms. Cost-sensitive CBO would simply include optimizable contextual variables as part of the design variables based on their cost. Instead, we adaptively select a subset of contextual variables to include in the optimization, based on the trade-off between their relevance and the additional cost incurred by optimizing them compared to leaving them to be determined by the environment. We learn the relevance of contextual variables by sensitivity analysis of the posterior surrogate model while minimizing the cost of optimization by leveraging recent developments on early stopping for BO. We empirically evaluate our proposed Sensitivity-Analysis-Driven Contextual BO (SADCBO) method against alternatives on both synthetic and real-world experiments, together with extensive ablation studies, and demonstrate a consistent improvement across examples.
@@ 1 Introduction
Bayesian optimization (BO) is a sample-efficient black-box optimization method, typically used when the objective function is too expensive to optimize directly (Garnett, 2023). Given an objective function that can be evaluated pointwise over a set of design variables, BO combines surrogate modeling with a pre-specified policy of evaluation over the design space (the so-called acquisition function) to efficiently locate the global optimum of the function. BO has been especially useful in automatic discovery of materials (Zhang et al., 2020), molecules (Fang et al., 2021), and pharmaceutical compounds (Gómez-Bombarelli et al., 2018; Korovina et al., 2020)—problem domains in which evaluating the performance of a candidate depends on a costly experiment.
@@
Despite the success of BO and recent algorithmic advancements, open challenges still remain for its practical use. A key implicit assumption in vanilla BO is that the objective function only depends on the design variables. This assumption is violated in many practical scenarios, wherein various environmental factors and experimental settings, referred to as contextual variables (Krause and Ong, 2011; Kirschner et al., 2020; Arsenyan et al., 2023), also affect the objective function. For instance, ambient humidity was found to influence the experiments in robot-assisted material design (Nega et al., 2021), such that the best process differed with humidity conditions. Moreover, in practice, the domain experts themselves might not know a priori which contextual variables are relevant, and would observe their confounding effect only during the course of the optimization process. Therefore, it is critical to identify the contextual variables that significantly affect the objective function: not only to guarantee reliable optimization results, but also for the practitioners to reliably reproduce experimental results.
@@
To deal with the uncertainty related to the contextual variables, variants of BO have been developed. In particular, Krause and Ong (2011) introduced the Contextual Bayesian optimization (CBO) framework, which uses the uncontrollable contextual information known before the experiment, like current environmental conditions, to enhance the surrogate model. Alternatively, several works have proposed to alter the simple optimization objective to make it robust in some sense, such as by taking the expectation w.r.t. the contextual variables (Toscano-Palmerin and Frazier, 2022), or considering distributionally-robust scenarios (Bogunovic et al., 2018; Kirschner et al., 2020). However, these works consider a different setup than the original CBO framework, as contextual information is only revealed after the design has been sent for experiment, not before. Besides this distinction, in some applications, contextual variables can be controlled, and therefore set to a value that was otherwise unlikely to be observed. For instance, synthesis conditions of material samples, such as sintering temperature or the used solvents, or certain environment conditions, such as experiment room temperature or ambient humidity (Higgins et al., 2021; Nega et al., 2021), are principally controllable during the course of an experiment, but it may not be straightforward to predict whether their inclusion is relevant (Abolhasani and Brown, 2023). Optimizing over all the potentially relevant contextual variables can improve BO performance, but this process can be costly, thus invoking a cost-vs.-efficiency trade-off.
Contributions.
@@
In this paper, we extend the CBO framework to settings in which the relevance of contextual variables is (i) not known beforehand, and (ii) can be optimized but at some cost. We propose a Sensitivity-Analysis-Driven CBO (SADCBO) algorithm for the simultaneous identification and optimization of relevant contextual variables. SADCBO leverages recent advances in sensitivity-analysis-driven variable selection (Sebenius et al., 2022) and early stopping criteria for BO (Ishibashi et al., 2023). We emphasize that SADCBO combines the contextual observational setting, where the context information is only observed, and the contextual optimization setting, where contextual variables can be optimized (similar to design variables), into a sequential algorithm. In effect, SADCBO provides a way to navigate the following tradeoff: should contextual variables be taken as is at no cost, or should they be steered outside of their observational distribution in order to provide more information about the objective, at a cost. We thoroughly evaluate the performance of SADCBO, comparing against methods from the CBO and high-dimensional BO literature, on both synthetic and real-world examples. SADCBO is shown to compare favorably to existing methods.
@@ 2 Contextual Bayesian Optimization (CBO)
The CBO framework (Krause and Ong, 2011) deals with a black-box function $f:\mathcal{X}\times\mathcal{Z}\rightarrow\mathbb{R}$ defined on the joint space of both the design variables $\mathcal{X}\subset\mathbb{R}^{d}$ and contextual variables $\mathcal{Z}\subset\mathbb{R}^{c}$. We assume that we get noisy evaluations of $f$, that is, we observe the output $y=f(\mathbf{x},\mathbf{z})+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma^{2}_{\text{noise}})$. A Gaussian process (GP) prior (Rasmussen and Williams, 2006) is placed on $f$; with the notation $\mathbf{v}=[\mathbf{x},\mathbf{z}]$, we write $f(\mathbf{v})\sim\mathcal{GP}(0,k(\mathbf{v},\mathbf{v}^{\prime}))$. A GP is a stochastic process fully characterized by its mean function (taken here to be zero without loss of generality) and its kernel $k(\mathbf{v},\mathbf{v}^{\prime})=\text{cov}[f(\mathbf{v}),f(\mathbf{v}^{\prime})]$. For any finite-dimensional collection of inputs $[\mathbf{v}_{1},\dots,\mathbf{v}_{t}]$, the function values $\mathbf{f}=[f(\mathbf{v}_{1}),\dots,f(\mathbf{v}_{t})]^{\top}\in\mathbb{R}^{t}$ follow a multivariate normal distribution $\mathbf{f}\sim\mathcal{N}(\mathbf{0},\mathbf{K})$, where $\mathbf{K}=(k(\mathbf{v}_{i},\mathbf{v}_{j}))_{1\leq i,j\leq t}$ is the kernel matrix. Given a dataset $\mathcal{D}_{t}=\{(\mathbf{x}_{i},\mathbf{z}_{i},y_{i})\}_{i=1}^{t}=\{(\mathbf{v}_{i},y_{i})\}_{i=1}^{t}$, the posterior distribution of $f(\mathbf{v})$ given $\mathcal{D}_{t}$ is Gaussian, with analytical expressions for the mean $\mu_{t}(\mathbf{v}|\mathcal{D}_{t})$ and variance $\sigma_{t}^{2}(\mathbf{v}|\mathcal{D}_{t})$.
@@
Given $\mathbf{z}_{t+1}$ and the previous $t$ observations $\mathcal{D}_{t}$, the next candidate design point $\mathbf{x}_{t+1}$ can be selected using the Upper Confidence Bound (UCB) acquisition function (Srinivas et al., 2012):
$\mathbf{x}_{t+1}$ $=\underset{\mathbf{x}\in\mathcal{X}}{\operatorname*{arg\,max}}~{}\alpha(\mathbf{x},\mathbf{z}_{t+1}|\mathcal{D}_{t})$
@@
We extend the problem setting of CBO in two ways. Firstly, we assume that only a subset of the contextual variables truly affect $f$. Let $\mathbf{z}=[z^{(1)},\dots,z^{(c)}]$ be the vector of all contextual variables. For any set $J$ belonging to the power set of $\{1,\dots,c\}$, denote by $\mathbf{z}^{(J)}\in\mathbb{R}^{|J|}$ the vector of reduced dimension whose variables are indexed by $J$. For instance, if $J=\{1,3\}$, then $\mathbf{z}^{(J)}=[z^{(1)},z^{(3)}]$. We assume that there exists a set $J^{\star}$, where $|J^{\star}|\ll c$, such that $f(\mathbf{x},\mathbf{z})=f(\mathbf{x},\mathbf{z}^{(J^{\star})})~{}\forall(\mathbf{x},\mathbf{z})$. Secondly, we include the possibility to set the value of any of the contextual variables at some cost, additionally to the usual design query cost $\lambda_{\mathbf{x}}$. This means that for all $j\in\{1,\dots,c\}$, the context variable $z^{(j)}$ can be optimized at a cost $\lambda_{j}$. To be able to control each contextual variable, we must also assume that they are independent: $p(\mathbf{z})=\prod_{j=1}^{c}p(z^{(j)})$. With these additional assumptions, we aim to maximize the function $f$ in a cost-efficient manner, while identifying the optimal set $J^{\star}$. This provides the user with a comprehensive summary of the relevant contextual variables found through optimization, thus ensuring reproducibility and explainability. Unlike CBO, the ability to control contextual variables allows us to judge whether or not one should optimize contextual variables to learn more about the objective (albeit at a cost), or if the current sampled context is already informative enough. Specifically, we aim to maximize the objective
$(\mathbf{x}_{t+1}^{\star},\mathbf{z}_{t+1}^{\star})=\operatorname*{arg\,max}_{(\mathbf{x},\mathbf{z}^{(J^{\star})}_{t+1})\in\mathcal{X}\times\prod_{j\in J^{\star}}\mathcal{Z}_{j}}~{}f(\mathbf{x},\mathbf{z}_{t+1})$ (3)
@@ 3 Methodology
To solve the extended CBO problem introduced in Section 2, we identify relevant contextual variables, building on a variable selection technique from the GP literature (Sebenius et al., 2022). Section 3.1 describes our adaption of this method to the optimization setting, by restricting the dataset to high function values. Section 3.2 then presents our sequential algorithm SADCBO, which employs the adapted variable selection method in solving the optimization problem.
@@ 3.1 Variable selection for CBO via sensitivity analysis
To handle the presence of contextual variables that can be optimized, one approach is to include them in the design space. However, such a strategy can be infeasible when their relevance is not known a priori and domain experts can only provide a candidate set of potentially relevant contextual variables, leading to an exponential expansion of the search space, or when the cost of optimizing all context variables is too high. In such cases, it is crucial to identify the relevant contextual variables, i.e. to find (a good approximation to) the optimal set $J^{\star}$. This not only allows us to optimize the function more efficiently, but also provides additional insights about the experiment to the domain experts.
To approximate the optimal set $J^{\star}$, we include those contextual variables that are most relevant towards identifying the optimum, which we estimate using sensitivity analysis. Specifically, we adapt the Feature Collapsing (FC) method (Sebenius et al., 2022). The FC method perturbs training points (namely, by setting one feature to zero), and measures the induced shift in the posterior predictive distribution in terms of KL divergence. Given a dataset $\mathcal{D}_{t}=\{(\mathbf{x}_{i},\mathbf{z}_{i},y_{i})\}_{i=1}^{t}$, the relevance $r_{i,j}$ of the $i^{\textup{th}}$ sample of the $j^{\textup{th}}$ contextual variable $z_{i}^{(j)}$ is computed as
@@
where $\boldsymbol{\xi}[j]=[\xi^{(1)},\dots,\xi^{(c)}]$ is a vector so that $\xi^{(j)}=0$, and $\xi^{(j^{\prime})}=1$ for $j^{\prime}\neq j$, and $\odot$ is the element-wise multiplication. The relevance score of the $j^{\textup{th}}$ contextual variable is then computed as
@@
The FC scores computed in this manner reveal the variables that are relevant for predicting the output across $\mathcal{D}_{t}$. As our goal is to maximize $f$, we are interested in identifying contextual variables that are relevant for high function values. Hence, we adapt Equation 5 to the BO setting by modifying the dataset over which the scores are averaged. We use information about high function values from two different sets: (1) The subset $\mathcal{D}^{\gamma_{t}}$ of the highest output values observed so far:
$\mathcal{D}^{\gamma_{t}}_{t}=\{(\mathbf{x}_{i},\mathbf{z}_{i},y_{i})\in\mathcal{D}_{t}~{}|~{}y_{i}/y_{\text{best}}\geq\gamma_{t}\},$ (6)
where $y_{\text{best}}=\max_{1\leq i\leq t}~{}y_{i}$ is the current observed maximum. For example, using $\gamma_{t}=0.8~{}\forall t$ would yield a $\mathcal{D}^{\gamma_{t}}_{t}$ that consists of the highest $20\%$ observations so far. (2) We select a batch of $Q$ points $\mathcal{D}^{Q}_{t}:=\{(\mathbf{x}^{\star}_{q},\mathbf{z}_{t+1})\}_{q=1}^{Q}$ that are promising for the next context $\mathbf{z}_{t+1}$:
$\{\mathbf{x}^{\star}_{q}\}_{q=1}^{Q}=\underset{\{\mathbf{x}_{q}\}_{q=1}^{Q}\in\mathcal{X}^{Q}}{\operatorname*{arg\,max}}~{}\alpha^{\text{Batch}}(\{(\mathbf{x}_{q},\mathbf{z}_{t+1})\}_{q=1}^{Q}|\mathcal{D}_{t}),$ (7)
where $\alpha^{\text{Batch}}$ denotes a batched version of the acquisition function $\alpha$ such as $Q$-UCB for UCB (Wilson et al., 2017). We use the union $\mathcal{D}_{t}^{\text{BO}}=\mathcal{D}^{\gamma_{t}}_{t}\cup\mathcal{D}^{Q}_{t}$ as our dataset for FC. Therefore, we compute $\text{FC}_{\mathcal{D}_{t}^{\text{BO}}}$ based on Equation 5. The importance of working with $\mathcal{D}_{t}^{\text{BO}}$ instead of $\mathcal{D}_{t}$ is illustrated in Figure 1 on a toy example.
We successively select the indices of the contextual variables with the highest FC scores until their cumulative FC score exceeds some chosen threshold $\eta\in[0,1]$, meaning that the selected variables explain the fraction $\eta$ of the output sensitivity amongst all contextual variables. Let $J_{\eta}$ denote the set of indices of the selected contextual variables. We train a GP surrogate based on $\{(\mathbf{x}_{i},\mathbf{z}_{i}^{(J_{\eta})},\mathbf{y}_{i})\}_{i=1}^{t}$ and can select a new design through maximization of the acquisition function $\alpha$:
$\mathbf{x}_{t+1}=\underset{\mathbf{x}\in\mathcal{X}}{\operatorname*{arg\,max}}~{}\alpha(\mathbf{x},\mathbf{z}^{(J_{\eta})}_{t+1}|\mathcal{D}_{t}).$ (8)
Note that other measures of variable relevance could have been used instead, e.g. the method proposed by Spagnol et al. (2019) based on maximum mean discrepancy (Gretton et al., 2012). We found FC to perform better, see Section 5.2.

### v4

@@ Abstract
Contextual Bayesian Optimization (CBO) efficiently optimizes black-box functions with respect to design variables, while simultaneously integrating contextual information regarding the environment, such as experimental conditions. However, the relevance of contextual variables is not necessarily known beforehand. Moreover, contextual variables can sometimes be optimized themselves at an additional cost, a setting overlooked by current CBO algorithms. Cost-sensitive CBO would simply include optimizable contextual variables as part of the design variables based on their cost. Instead, we adaptively select a subset of contextual variables to include in the optimization, based on the trade-off between their relevance and the additional cost incurred by optimizing them compared to leaving them to be determined by the environment. We learn the relevance of contextual variables by sensitivity analysis of the posterior surrogate model while minimizing the cost of optimization by leveraging recent developments on early stopping for BO. We empirically evaluate our proposed Sensitivity-Analysis-Driven Contextual BO (SADCBO) method against alternatives on both synthetic and real-world experiments, together with extensive ablation studies, and demonstrate a consistent improvement across examples.
@@ 1 Introduction
Bayesian optimization (BO) is a sample-efficient black-box optimization method, typically used when the objective function is too expensive to optimize directly [Garnett, 2023]. Given an objective function that can be evaluated pointwise over a set of design variables, BO combines surrogate modeling with a pre-specified policy of evaluation over the design space (the so-called acquisition function) to efficiently locate the global optimum of the function. BO has been especially useful in automatic discovery of materials [Zhang et al., 2020], molecules [Fang et al., 2021], and pharmaceutical compounds [Gómez-Bombarelli et al., 2018, Korovina et al., 2020]—problem domains in which evaluating the performance of a candidate depends on a costly experiment.
@@
Despite the success of BO and its recent algorithmic advancements, open challenges remain for its practical use. A key implicit assumption in vanilla BO is that the objective function only depends on the design variables. This assumption is violated in many practical scenarios, wherein various uncontrolled environmental factors and experimental settings, referred to as contextual variables [Krause and Ong, 2011, Kirschner et al., 2020, Arsenyan et al., 2023], also affect the objective function. For instance, ambient humidity was found to influence the experiments in robot-assisted material design [Nega et al., 2021], such that the best compound differed with humidity conditions. Moreover, in practice, the domain experts themselves might not know a priori which contextual variables are relevant, and would observe their confounding effect only during the course of the optimization process. Therefore, it is critical to identify the contextual variables that significantly affect the objective function, not only to achieve the highest optimization results, but also for the practitioners to reliably reproduce experimental results.
@@
To deal with the uncertainty related to the contextual variables, variants of BO have been developed. In particular, Krause and Ong [2011] introduced the Contextual Bayesian optimization (CBO) framework, which uses the uncontrollable contextual information known before the experiment, like current environmental conditions, to enhance the surrogate model. Alternatively, several works have proposed to alter the simple optimization objective to make it robust in some sense, such as by taking the expectation with respect to the contextual variables [Toscano-Palmerin and Frazier, 2022], or considering distributionally-robust scenarios [Bogunovic et al., 2018, Kirschner et al., 2020]. However, these works consider a different setup than the original CBO framework, as contextual information is only revealed after the design has been sent for experiment, not before. Besides this distinction, in some applications, contextual variables can be controlled, and therefore set to values they may be unlikely to take during passing observation. Such variables are, for instance, synthesis conditions of material samples, including sintering temperature or the used solvents. Certain environmental conditions like room temperature or ambient humidity are also “principally” controllable during the course of an experiment [Higgins et al., 2021, Nega et al., 2021]. Nevertheless, whether their inclusion as optimization variables is relevant or not may not be straightforward to predict [Abolhasani and Brown, 2023]. Moreover, optimizing over all the potentially relevant contextual variables can improve BO performance, but this process can be costly, thus invoking a cost-versus-efficiency trade-off.
@@
Contributions. In this paper, we extend the CBO framework to settings in which the relevance of contextual variables is (i) not known beforehand, and (ii) can be optimized, but at some cost. We propose a Sensitivity-Analysis-Driven CBO (SADCBO) algorithm for the simultaneous identification and optimization of relevant contextual variables. SADCBO leverages recent advances in sensitivity-analysis-driven variable selection [Sebenius et al., 2022] and early stopping criteria for BO [Ishibashi et al., 2023]. We emphasize that SADCBO combines the contextual observational setting, where the context information is only observed, and the contextual optimization setting, where contextual variables can be optimized (similar to design variables), into a sequential algorithm. In effect, SADCBO provides a way to navigate the following tradeoff: should contextual variables be taken as is at no cost, or should they be steered outside of their observational distribution in order to provide more information about the objective, at a cost? We evaluate the performance of SADCBO, comparing against methods from the CBO and high-dimensional BO literature, on both synthetic and real-world cases.
@@ 2 Contextual Bayesian Optimization (CBO)
The CBO framework [Krause and Ong, 2011] deals with a black-box function $f:\mathcal{X}\times\mathcal{Z}\rightarrow\mathbb{R}$ defined on the joint space of both the design variables $\mathcal{X}\subset\mathbb{R}^{d}$ and contextual variables $\mathcal{Z}\subset\mathbb{R}^{c}$. We assume that we get noisy evaluations of $f$, that is, we observe the output $y=f(\mathbf{x},\mathbf{z})+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma^{2}_{\text{noise}})$. A Gaussian process (GP) prior [Rasmussen and Williams, 2006] is placed on $f$; with the notation $\mathbf{v}=[\mathbf{x},\mathbf{z}]$, we write $f(\mathbf{v})\sim\mathcal{GP}(0,k(\mathbf{v},\mathbf{v}^{\prime}))$. A GP is a stochastic process fully characterized by its mean function (taken here to be zero) and its kernel $k(\mathbf{v},\mathbf{v}^{\prime})=\text{cov}[f(\mathbf{v}),f(\mathbf{v}^{\prime})]$. This implies that for any finite-dimensional collection of inputs $[\mathbf{v}_{1},\dots,\mathbf{v}_{t}]$, the function values $\mathbf{f}=[f(\mathbf{v}_{1}),\dots,f(\mathbf{v}_{t})]^{\top}\in\mathbb{R}^{t}$ follow a multivariate normal distribution $\mathbf{f}\sim\mathcal{N}(\mathbf{0},\mathbf{K})$, where $\mathbf{K}=(k(\mathbf{v}_{i},\mathbf{v}_{j}))_{1\leq i,j\leq t}$ is the kernel matrix. Given a dataset $\mathcal{D}_{t}=\{(\mathbf{x}_{i},\mathbf{z}_{i},y_{i})\}_{i=1}^{t}=\{(\mathbf{v}_{i},y_{i})\}_{i=1}^{t}$, the posterior distribution of $f(\mathbf{v})$ given $\mathcal{D}_{t}$ is Gaussian, with analytical expressions for the mean $\mu_{t}(\mathbf{v}|\mathcal{D}_{t})$ and variance $\sigma_{t}^{2}(\mathbf{v}|\mathcal{D}_{t})$.
@@
Given $\mathbf{z}_{t+1}$ and the previous $t$ observations $\mathcal{D}_{t}$, the next candidate design $\mathbf{x}_{t+1}$ is selected using the Upper Confidence Bound (UCB) acquisition function $\alpha$ [Srinivas et al., 2012]:
$\mathbf{x}_{t+1}$ $=\underset{\mathbf{x}\in\mathcal{X}}{\operatorname*{arg\,max}}\leavevmode\nobreak\ \alpha(\mathbf{x},\mathbf{z}_{t+1}|\mathcal{D}_{t})$
@@
We extend the problem setting of CBO in two ways. Firstly, we assume that only a subset of the contextual variables truly affect $f$. Let $\mathbf{z}=[z^{(1)},\dots,z^{(c)}]$ be the vector of all contextual variables. For any set $J$ belonging to the power set of $\{1,\dots,c\}$, denote by $\mathbf{z}^{(J)}\in\mathbb{R}^{|J|}$ the vector of reduced dimension whose variables are indexed by $J$. For instance, if $J=\{1,3\}$, then $\mathbf{z}^{(J)}=[z^{(1)},z^{(3)}]$. We assume that there exists a set $J^{\star}$, where $|J^{\star}|\ll c$, such that $f(\mathbf{x},\mathbf{z})=f(\mathbf{x},\mathbf{z}^{(J^{\star})})\leavevmode\nobreak\ \forall(\mathbf{x},\mathbf{z})$. Secondly, we include the possibility of setting the value of any of the contextual variables at some cost over and above the usual design query cost $\lambda_{\mathbf{x}}$. This means that for all $j\in\{1,\dots,c\}$, the context variable $z^{(j)}$ can be optimized at a cost $\lambda_{j}$. To be able to control each contextual variable, we must also assume their independence: $p(\mathbf{z})=\prod_{j=1}^{c}p(z^{(j)})$. With these additional assumptions, we aim to maximize the function $f$ in a cost-efficient manner, while identifying the optimal set $J^{\star}$. This provides the user with a comprehensive summary of the relevant contextual variables found through optimization, thus ensuring reproducibility and explainability. Unlike CBO, the ability to control contextual variables allows us to judge whether or not one should optimize contextual variables to learn more about the objective (albeit at a cost), or if the current sampled context is already informative enough. Specifically, we aim to maximize the objective
$(\mathbf{x}_{t+1}^{\star},\mathbf{z}_{t+1}^{\star})=\operatorname*{arg\,max}_{(\mathbf{x},\mathbf{z}^{(J^{\star})}_{t+1})\in\mathcal{X}\times\prod_{j\in J^{\star}}\mathcal{Z}_{j}}\leavevmode\nobreak\ f(\mathbf{x},\mathbf{z}_{t+1})$ (3)
@@ 3 Methodology
To solve the extended CBO problem introduced in Section 2, we identify relevant contextual variables, building on a variable selection technique from the GP literature [Sebenius et al., 2022]. Section 3.1 describes our adaptation of this method to the optimization setting, by restricting the dataset to high function values. Section 3.2 then presents our sequential algorithm SADCBO, which employs the adapted variable selection method in solving the optimization problem. A flowchart summarizing the proposed method can be found in Figure S1.
@@ 3.1 Variable selection for CBO via sensitivity analysis
To handle the presence of contextual variables that can be optimized, one approach is to include them in the design space. However, such a strategy can be infeasible when their relevance is not known a priori and domain experts can only provide a candidate set of potentially relevant contextual variables. Indeed, this leads to an exponential expansion of the search space, while at the same time increasing the cost of optimization. In such cases, it is crucial to identify the relevant contextual variables, i.e., to find (a good approximation to) the optimal set $J^{\star}$. This not only allows us to optimize the function more efficiently but also provides additional insights about the experiment to the domain experts.
To approximate the optimal set $J^{\star}$, we include those contextual variables that are most relevant for identifying the optimum, which we estimate using sensitivity analysis. Specifically, we adapt the Feature Collapsing (FC) method [Sebenius et al., 2022]. The FC method perturbs training points (namely, by setting one feature to zero), and measures the induced shift in the posterior predictive distribution in terms of KL divergence. Given a dataset $\mathcal{D}_{t}=\{(\mathbf{x}_{i},\mathbf{z}_{i},y_{i})\}_{i=1}^{t}$, the relevance $r_{i,j}$ on the $i^{\textup{th}}$ sample of the $j^{\textup{th}}$ contextual variable $z_{i}^{(j)}$ is computed as
@@
where $\boldsymbol{\xi}[j]=[\xi^{(1)},\dots,\xi^{(c)}]$ is a vector so that $\xi^{(j)}=0$, and $\xi^{(j^{\prime})}=1$ for $j^{\prime}\neq j$, and $\odot$ is the element-wise multiplication. The relevance score of the $j^{\textup{th}}$ contextual variable is then computed as an average over $\mathcal{D}_{t}$:
@@
The FC scores obtained in this manner reveal the variables that are relevant for predicting the output across $\mathcal{D}_{t}$. As our goal is to maximize $f$, we are interested in identifying contextual variables that are relevant for high function values. Hence, we adapt Equation 5 to the BO setting by modifying the dataset over which the scores are averaged. We use information about high function values from two different sets: (1) The subset $\mathcal{D}^{\gamma_{t}}$ associated with the highest output values observed so far:
$\mathcal{D}^{\gamma_{t}}_{t}=\{(\mathbf{x}_{i},\mathbf{z}_{i},y_{i})\in\mathcal{D}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ y_{i}/y_{\text{best}}\geq\gamma_{t}\},$ (6)
where $y_{\text{best}}=\max_{1\leq i\leq t}\leavevmode\nobreak\ y_{i}$ is the current observed maximum. For example, using $\gamma_{t}=0.8\leavevmode\nobreak\ \forall t$ would yield a $\mathcal{D}^{\gamma_{t}}_{t}$ that consists of the highest $20\%$ observations so far. (2) We select a batch of $Q$ points $\mathcal{D}^{Q}_{t}:=\{(\mathbf{x}^{\star}_{q},\mathbf{z}_{t+1})\}_{q=1}^{Q}$ that are promising given the next context $\mathbf{z}_{t+1}$:
$\{\mathbf{x}^{\star}_{q}\}_{q=1}^{Q}=\underset{\{\mathbf{x}_{q}\}_{q=1}^{Q}\in\mathcal{X}^{Q}}{\operatorname*{arg\,max}}\leavevmode\nobreak\ \alpha^{\text{Batch}}(\{(\mathbf{x}_{q},\mathbf{z}_{t+1})\}_{q=1}^{Q}|\mathcal{D}_{t}),$ (7)
where $\alpha^{\text{Batch}}$ denotes a batched version of the acquisition function $\alpha$ such as $Q$-UCB for UCB [Wilson et al., 2017]. We use the union $\mathcal{D}_{t}^{\text{BO}}=\mathcal{D}^{\gamma_{t}}_{t}\cup\mathcal{D}^{Q}_{t}$ as our dataset for FC. Therefore, we compute $\text{FC}_{\mathcal{D}_{t}^{\text{BO}}}$ based on Equation 5. The importance of working with $\mathcal{D}_{t}^{\text{BO}}$ instead of $\mathcal{D}_{t}$ is illustrated in Figure 1 on a toy example.
We successively select the indices of the contextual variables with the highest FC scores until their cumulative FC score exceeds some chosen threshold $\eta\in[0,1]$, meaning that the selected variables explain the fraction $\eta$ of the output sensitivity amongst all contextual variables. Let $J_{\eta}$ denote the set of indices of the selected contextual variables. We train a GP surrogate based on $\{(\mathbf{x}_{i},\mathbf{z}_{i}^{(J_{\eta})},y_{i})\}_{i=1}^{t}$ and can select a new design through maximization of the acquisition function $\alpha$:
$\mathbf{x}_{t+1}=\underset{\mathbf{x}\in\mathcal{X}}{\operatorname*{arg\,max}}\leavevmode\nobreak\ \alpha(\mathbf{x},\mathbf{z}^{(J_{\eta})}_{t+1}|\mathcal{D}_{t}).$ (8)
Note that other measures of variable relevance could have been used, e.g., the method proposed by Spagnol et al. [2019] based on maximum mean discrepancy [Gretton et al., 2012]. We found FC to perform better (see Section 5.2).